OpenStudy (anonymous):

Brian has reduced his cholesterol level by 7% by following a strict diet and regular exercise. If his original level was 300, what is his cholesterol level now? 21 298 321 279

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

i think it is 379

OpenStudy (wolfe8):

If it's less 7% then he has only 93% of the original now. You cannot have more than before if you 'reduced'.

OpenStudy (anonymous):

21

OpenStudy (wolfe8):

93% of 300 can be translated into 93/100 * 300

OpenStudy (anonymous):

sorry i ment 279

OpenStudy (anonymous):

321

OpenStudy (wolfe8):

Did you even do the equation I put out?

OpenStudy (anonymous):

@Chris911

OpenStudy (anonymous):

@Eli_Moses

OpenStudy (anonymous):

@LieutenantGeneral

OpenStudy (anonymous):

@praxer

OpenStudy (praxer):

It will be original minus the rate of decrease * the original value = 300- 7%300 =300-7/100*300 =300-21 =279 The option D. is correct....

OpenStudy (anonymous):

preciate it

OpenStudy (praxer):

$$\huge \color{red}{\ddot\smile}......$$